# E rror S pan A nnotation: A Balanced Approach for Human Evaluation of Machine Translation

### v1

@@ Abstract
High-quality Machine Translation (MT) evaluation relies heavily on human judgments. Comprehensive error classification methods, such as Multidimensional Quality Metrics (MQM), are expensive as they are time-consuming and can only be done by experts, whose availability may be limited especially for low-resource languages. On the other hand, just assigning overall scores, like Direct Assessment (DA), is simpler and faster and can be done by translators of any level, but are less reliable. In this paper, we introduce Error Span Annotation (ESA), a human evaluation protocol which combines the continuous rating of DA with the high-level error severity span marking of MQM. We validate ESA by comparing it to MQM and DA for 12 MT systems and one human reference translation (English to German) from WMT23. The results show that ESA offers faster and cheaper annotations than MQM at the same quality level, without the requirement of expensive MQM experts.
footnotetext: Equal contributions. Others alphabetically.footnotetext: 0Code & data will be released before WMT 2024 at:
github.com/wmt-conference/ErrorSpanAnnotations
@@ 1 Introduction
A simpler alternative in the form of highlighting errors without assigning their error types is a compromise between overall scoring and error classification Kreutzer et al. (2020); Popović (2020). While the previously reported findings on this method are promising, no systematic comparison for the purposes of evaluating machine translation systems has been carried out so far.
@@ 2 Related work
Several error span marking methods have been proposed recently Kreutzer et al. (2020); Popović (2020) as a less demanding error annotation approach than error classification. While it does not provide the fine-grained details about different error classes, it still gives the information about the position and amount of errors, and also enables further fine-grained analysis on the annotated data, if necessary (e.g. classification of already marked errors, identifying linguistic phenomena causing the errors, or focusing on particular error type). Although promising, the potential and limits of error marking have not been systematically investigated so far. Specifically, they were not directly compared to other methods and used in the context of WMT campaign evaluation and system ranking.
3 Comparison DA+SQM and MQM
@@
As an example, in App. Figure 8 we show the system ranking based on two approaches DA+SQM and MQM on Chinese-English (sentence-level evaluation) and English-German (paragraph-level evaluation) language pairs. Although both technique reach same order of system clusters, DA+SQM produces much fewer clusters in paragraph-level setup, thus putting many systems within a single cluster. On the other hand, MQM is better able to distinguish different systems. To increase statistical power of DA+SQM, we would have to collect much more DA+SQM samples (Wei et al., 2022), which would further drive the cost up. In addition, DA is much more skewed towards fluency as opposed to adequacy (Martindale and Carpuat, 2018).
@@ 4 Error Span Annotation
In Error Span Annotation (ESA), the evaluators first mark all erroneous or problematic parts (characters, words, phrases, sentences) in the translated text. For each marked span, they are asked to provide one of the two severity levels: major (e.g. changed meaning) or minor (e.g. incorrect grammar, style, etc.; see Table 1). Because all error span annotations are made in the translation, not the source text, we include a special tag for marking omission errors. This was an intentional design choice over annotating the source text to make the annotation protocol forward-compatible with other translation modalities, such as audio and video translations.
After the annotators mark all the error spans, they provide an overall score for the entire segment, on the scale from 0 to 100, reminiscent of DA+SQM. We implement this protocol annotation interface in Appraise (Federmann, 2018) and show a screenshot in Footnote 2. The full guidelines displayed to annotators are shown in Appendix A.
@@
To rank systems, we need scalar values. There are two evident ways to extract them from ESA: (1) using the annotator’s overall segment-level score directly, like DA+SQM, or (2) converting error span severity levels into a segment-level score, like MQM. By instructing the annotators to identify and mark all errors first, we prime them to be more accurate when assessing the overall quality of the segment—when making the decision about the score, they see all marked errors in the segment and can take them into consideration.
MQM is primarily error diagnostics protocol which has been repurposed for translation segment scoring. The transition from error spans into a single segment score has been proposed to be done with the formula based on error severity counts (Freitag et al., 2021a): [^3] The MQM score calculation weights “untranslated” with -25 and “fluency/punctuation” with -0.1. ESA does not follow this due to lacking error categories.
@@
Notice that this does not scale with different text sizes. As an example, if translation of a segment has two major errors, it receives the score of $-10$. However, if the source is repeated twice and the corresponding translation as well, the score would be further decreased to $-20$. This is especially problematic for paragraph-level evaluation which features segments of different length. Additionally, the segment-level MQM score might not correspond with the segment-level translation quality, such as when marking one error affecting several places in the segment, as in Table 1 (top). To avoid such issues, we use the annotators’ direct scores as the main scoring approach for the ESA protocol unless specified differently. The system-level scores for all human evaluation methods in this work are calculated as the average of all segment-level scores for particular system. We further revisit the score computation in App. Section B.1. In some analysis, we use ESA error spans to calculate MQM-like score, we refer to such scores as ESAspans.
@@ 5 Experimental setup
We conduct experiments comparing the ESA protocol with MQM and DA+SQM protocols. We design them in a way, which makes it comparable with the previously collected annotations for MQMWMT and DA+SQMWMT. For this reason, we reproduce the human evaluation campaign for the WMT23 English to German systems (Kocmi et al., 2023; Freitag et al., 2023) with ESA and our reimplementation of MQM.
The original campaign featured 13 translations of 557 source segments. To facilitate running multiple campaigns for proper comparison, we had to scale down and subsampled 207 segments per system (74 documents), which yields 2,691 segments in total. To keep the ESA annotation comparable to other protocols, we subsample by selecting a subset of documents evaluated by Freitag et al. (2023) keeping the entire documents. This differs from Freitag et al. (2023), who removed some paragraphs from the ends of long documents. In our analysis we only consider segments overlapping with both previous annotation collections, thus obtaining 2,027 annotations evaluated across all protocols. We note that the subsampling makes all annotations protocols statistically less powerful, but keeps them fair in terms of statistical power per evaluated segment. Therefore, clustering and final system ranking in our analysis differs from Kocmi et al. (2023); Freitag et al. (2023). To keep the study comparable with Kocmi et al. (2023), we use the Wilcoxon rank-sum test with $p{<}0.05$ when producing system clusters. However, as all systems are evaluated on the same set of segments, we advice to use Wilcoxon signed-rank test when evaluation ESA as proposed by Kocmi et al. (2021).
@@
To analyze inter annotator agreement, we run ESA protocol twice with different sets of annotators. We hired 28 annotators to evaluate our protocols and each protocol was evaluated by different sets of experts for less biased results. Specifically, we had 8 bilingual annotators for the initial run of ESA1, 10 translators for ESA2 (different vendor), and 10 MQM professionals to evaluate MQM protocol. For MQM, we hired professionals already experienced with MQM annotation protocol, while for ESA, we hired translators or bilingual speakers. All of them were native speakers of the target language, i.e., German.
@@ 6.1 Score distribution
As per Table 2, on average the ESA1 annotators mark 0.45 error spans per segments, which is close to MQM’s 0.53 error spans per segment. The second run of ESA2 has more than double of error spans per segment, which could be the result of different characteristics of annotators group, specifically experienced annotators in ESA1 versus translators in ESA2. On the other hand, MQMWMT contains 7x more errors per segment than our rerun of MQM. The severity levels are distributed similarly across campaigns. Important insights are in the score range distribution presented in Figure 2: the MQM-like score computation creates more skewed distribution around 0, which is in addition unbounded and can go to -infinity the longer the evaluated segment is, complicating modeling and comparisons. In contrast, the manual scores from annotators are more spread out and guaranteed to be in the interval [0, 100].
@@ 6.2 System ranking capabilities
We now investigate how well different techniques can rank MT systems. For purpose of this experiment, we consider MQMWMT as the gold standard.
@@
The Figure 3 also shows, that relying on error spans only isn’t optimal, as ESAspans has lower accuracy and Spearman’s correlation than ESA. We can notice that this is even lower than our rerun of MQM. This can be attributed to the evaluation crowd, where we used professional MQM annotators for MQM protocol, while we used bilingual speakers and translators for the ESA protocol.
@@ 6.3 Agreement with other protocols
We now compare how different protocols correspond on the segment-level. We use MQMWMT as the gold standard to compare against because it was done outside of our setup. We analyze two aspects: scores and spans. For this evaluation, we use Kendall $\tau$ variant C, which is more suitable for data with different underlying scales and many ties.
In Table 3, we see that although all protocols correlate similarly with MQMWMT, our protocols obtain the highest $\tau$ for both runs. The segment-level correlation also confirms that relying only on the error spans isn’t optimal and ESAspans obtains lower Kendall score.
@@
Table 4: Similarity between spans of different annotation protocols (one campaign) computed as percentage how much of $B$ does $A$ contain. For example, 93% of ESA spans were also in MQM. Any span that overlaps with another one is considered a hit.
@@ 6.4 Quality of annotations
Intra annotator agreement
@@
Lastly, evaluating inter annotator agreement is heavily affected by the strategy of annotators, where different annotation strategy doesn’t mean different performance of the task as Riley et al. (2024) showed. Secondly, the MQMWMT was collected with a different tooling and the documents have been presented to annotators in different order, which could also impact the inter annotator agreement.
@@
Inter annotator agreement
@@
Error span agreement.
@@ 6.5 Annotation time
One of the reasons behind development of the new protocol is reducing the time requirements of the evaluation. In this section, we analyze times on our experiments only, therefore our rerun of MQM in contrast to ESA. Assessing the speed of annotations is challenging as annotators took breaks during the annotation (from short breaks taking several minutes up to several hours). This makes the evaluation of time problematic and we therefore investigate the time estimate in several ways. Counting all annotators together, the median time for annotation per single paragraph for MQM is 38 seconds and for ESA is 29 seconds, a reduction of 23%. However, as median time for each annotator fluctuates, we look at the average median time across annotators. For MQM, the median is 49 seconds and for ESA it is 34 seconds, a reduction of 32%. This can be contributed mainly to the less demanding error span annotation approach used in ESA.
@@ 7 Conclusion
A possible limitation in contrast to MQM is that now the system evaluation does not provide breakdown of error types, which could help practitioners in improving their systems. ESA does not provide this, though the goal is primarily evaluation and not diagnosis. Because of this and the costs of scaling expert labor, we are convinced that this is not a true shortcoming of ESA. Furthermore, the annotated errors can be further classified and analysed, if necessary.
@@ Appendix A User Guidelines
The following are are annotation guidelines for our local ESA and MQM campaigns.
@@ A.2 MQM (Multidimensional Quality Metrics)
Appendix B Additional results
Figure 6: Intra annotator agreement - changes in scoring by the same annotator when evaluated again. Each point represents single annotated segment with x-axis being annotator’s score assigned in March and y-axis their score assigned in May.
B.1 From error spans to final score
@@
The MQM formula was crafted with respect to preserving system ranking and not segment-level matching from the same annotator (Freitag et al., 2021a). However, the construction of ESA allows us to revisit this problem as each annotator gives both the error spans and the final score. We scan for multiple minor/major ratios of error weights and show the results in Figure 7. We find that the optimal formula that optimizes the correlation between the direct score and the score from the spans has the following form:
[…]
which is very close to the originally proposed 1:5 ratio. From Figure 7, the weight of the major error class seems to have a much bigger effect on the final translation score, suggesting that minor errors play a lesser role.
Figure 9: Tutorial to ESA annotations shown at the beginning of the campaign. All segments needed to be annotated correctly before continuing.

### v2

@@ Abstract
High-quality Machine Translation (MT) evaluation relies heavily on human judgments. Comprehensive error classification methods, such as Multidimensional Quality Metrics (MQM), are expensive as they are time-consuming and can only be done by experts, whose availability may be limited especially for low-resource languages. On the other hand, just assigning overall scores, like Direct Assessment (DA), is simpler and faster and can be done by translators of any level, but is less reliable. In this paper, we introduce Error Span Annotation (ESA), a human evaluation protocol which combines the continuous rating of DA with the high-level error severity span marking of MQM. We validate ESA by comparing it to MQM and DA for 12 MT systems and one human reference translation (English to German) from WMT23. The results show that ESA offers faster and cheaper annotations than MQM at the same quality level, without the requirement of expensive MQM experts.
footnotetext: Equal contributions. Others alphabetically.footnotetext: 0Code & collected data:
[…]
@@ 2 Related work
Several error span marking methods have been proposed recently Kreutzer et al. (2020); Popović (2020) as a less demanding error annotation approach than error classification. While it does not provide the fine-grained details about different error classes, it still gives the information about the position and amount of errors, and also enables further fine-grained analysis on the annotated data, if necessary (e.g. classification of already marked errors, identifying linguistic phenomena causing the errors, or focusing on particular error type). While the previously reported findings on this method are promising, no systematic comparison for the purposes of evaluating machine translation systems has been carried out so far. Furthermore, the simplified error marking method does not solve the challenges in determining how to appropriately weight individual errors to obtain segment-level scores, a problem that becomes particularly pronounced when extending the evaluation to document level.
This work combines the advantages of error annotation (like MQM) and assigning direct scores (like DA). The annotators are first asked to identify and mark all errors, and afterwards to assign an overall score. When deciding about the score, they are primed by the preceding error annotation and see all the marked errors that can be taken into consideration for the final score.
3 Comparison with DA+SQM and MQM
@@
As an example, in App. Figure 7 we show the system ranking based on two approaches DA+SQM and MQM on Chinese-English (sentence-level evaluation) and English-German (paragraph-level evaluation) language pairs. Although both techniques reach same order of system clusters, DA+SQM produces much fewer clusters in paragraph-level setup, thus putting many systems within a single cluster. On the other hand, MQM is better able to distinguish different systems. To increase statistical power of DA+SQM, we would have to collect much more DA+SQM samples (Wei et al., 2022), which would further drive the cost up. In addition, DA is much more skewed towards fluency as opposed to adequacy (Martindale and Carpuat, 2018).
@@ 4 Error Span Annotation
In Error Span Annotation (ESA), the evaluators first mark all problematic parts (characters, words, phrases, sentences) in the translated text. For each marked span, they also provide one of the two severity levels: major (e.g. changed meaning) or minor (e.g. incorrect grammar, style; see Table 1). Because all error spans are marked in the translation, not the source text, we include a special tag for marking omission errors. This was an intentional choice over annotating the source text to make the annotation protocol forward-compatible with other translation modalities, such as audio and video translations.
After the annotators mark all the error spans, they are asked to provide an overall score for the entire segment, on the scale from 0 to 100, reminiscent of DA+SQM. We implement the annotation interface in Appraise (Federmann, 2018) and show a screenshot in Figure 2. The full guidelines displayed to annotators are shown in Appendix A.
@@
To rank systems, we need scalar values. There are two evident ways to extract them from ESA: (1) using the annotator’s overall segment-level score directly, like DA+SQM, or (2) converting error span severity levels into a segment-level score, like MQM. By instructing the annotators to identify and mark all errors first, we prime them to be more accurate when assessing the overall quality of the segment—when making the decision about the score, they have already marked all errors in the segment and can see them, therefore they can take them into consideration.
MQM is primarily error diagnostics protocol which has been repurposed for translation segment scoring. The transition from MQM error spans into a single segment score has been proposed to be done with the formula based on error severity counts (Freitag et al., 2021a):
@@
Notice that this does not scale with different text sizes. [^3] Another weighting was proposed by Burchardt (2013) that is less prevalent in machine translation evaluation. As an example, if translation of a segment has two major errors, it receives the score of $-10$. However, if the source is repeated twice and the corresponding translation as well, the score would be further decreased to $-20$. This is especially problematic for paragraph-level evaluation which features segments of different length. Additionally, the segment-level MQM score might not correspond with the segment-level translation quality, such as when marking one error affecting several places in the segment, as in Table 1 (top). To avoid such issues, we use the annotators’ direct scores as the main scoring approach for the ESA protocol unless specified differently. The system-level scores for all human evaluation methods in this work are calculated as the average of all segment-level scores for particular system. We further revisit the score computation in App. Section B.1. In some analysis, we use ESA error spans to calculate MQM-like score, we refer to such scores as ESAspans.
@@
Figure 2: Screenshot of the beginning of one annotated document in the ESA interface (following segments are not shown). By showing and annotating whole documents at the segment-level, the annotators see all the relevant context. Segment reset button and completed labels removed for brevity. See the interactive tutorial shown to all annotators in Appendix Figure 11.
@@ 5 Experimental setup
We conduct experiments comparing the ESA protocol with MQM and DA+SQM protocols. We design them in a way which makes it comparable with the previously collected annotations for MQMWMT and DA+SQMWMT. For this reason, we reproduce the human evaluation campaign for the WMT23 English to German systems (Kocmi et al., 2023; Freitag et al., 2023) with ESA and our reimplementation of MQM.
The original campaign featured 13 translations of 557 source segments. To facilitate running multiple campaigns for proper comparison, we had to scale down and subsampled 207 segments per system (74 documents), which yields 2,691 segments in total. To keep the ESA annotation comparable to other protocols, we subsample by selecting a subset of documents evaluated by Freitag et al. (2023) keeping the entire documents. This differs from Freitag et al. (2023), who removed some paragraphs from the ends of long documents. In our analysis we only consider segments overlapping with both previous annotation collections, thus obtaining 2,027 annotations evaluated across all protocols. We note that the subsampling makes all annotations protocols statistically less powerful, but keeps them fair in terms of statistical power per evaluated segment. Therefore, clustering and final system ranking in our analysis differs from Kocmi et al. (2023); Freitag et al. (2023). To keep the study comparable with Kocmi et al. (2023), we use the Wilcoxon rank-sum test with $p{<}0.05$ when producing system clusters. However, as all systems are evaluated on the same set of segments, we advice to use Wilcoxon signed-rank test when employing ESA as proposed by Kocmi et al. (2021).
@@
To analyze inter annotator agreement, we run ESA protocol twice with different sets of annotators. We hired 28 annotators to evaluate our protocols and each protocol was evaluated by different sets of experts to avoid bias. Specifically, we had 8 bilingual annotators for the initial run of ESA1, 10 translators for ESA2 (different vendor), and 10 annotators to evaluate MQM protocol. For MQM, we hired professionals already experienced with MQM annotation protocol, while for ESA, we hired translators or bilingual speakers. All of them were native speakers of the target language, German.
@@ 6 Analysis
We first analyze the collected data, system ranking, agreement with other protocols, quality assurance, and finally the annotation time. The findings reveal that the ESA quality is comparable, if not better than MQM, takes less time, and does not require highly trained annotators.
@@ 6.1 Score distribution
As per Table 2, on average the ESA1 annotators mark 0.45 error spans per segments, which is close to MQM’s 0.53 error spans per segment. The second run of ESA2 has more than double of error spans per segment, which could be the result of different characteristics of annotators group, specifically experienced annotators in ESA1 versus translators in ESA2. On the other hand, MQMWMT contains 7x more errors per segment than our re-run of MQM. We attribute this difference primarily to the differences in annotation crowds, which further motivates our own evaluation of both MQM and ESA so that the annotations differ only in the annotation protocol and past MQM training. The severity levels are distributed similarly across campaigns. Important insights are in the score range distribution presented in Figure 3: the MQM-like score computation creates more skewed distribution around 0, which is in addition unbounded and can go to -infinity the longer the evaluated segment is, complicating modeling and comparisons. In contrast, the manual scores from annotators are spread out and guaranteed to be in [0, 100].
@@ 6.2 System ranking capabilities
We now investigate how well different annotation protocols (ESA and MQM) can rank MT systems. For purpose of this experiment, we consider MQMWMT which comes from an independent very high-quality annotation crowd and implementation, as the gold standard. Note that this creates a positive bias towards our implementation of MQM as opposed to ESA. In Section B.2 we show an evaluation without this gold standard assumption.
@@
The Figure 4 also shows, that relying on error spans only is not optimal, as ESAspans has lower accuracy and Spearman’s correlation than ESA. We can notice that this is even lower than our rerun of MQM. This can be attributed to the evaluation crowd, where we used professional MQM annotators for MQM protocol, while we used bilingual speakers and translators for the ESA protocol.
@@ 6.3 Agreement with other protocols
We now compare how different protocols correspond on the segment-level. We use MQMWMT as the gold standard to compare against because it was done independently outside of our setup and with high quality assurance. We analyze two aspects: (1) segment scores, and (2) spans, where we consider spans overlapping even with a single character as a match. For this evaluation, we use Kendall $\tau$ variant C, which is more suitable for data with different underlying scales and many ties.
In Table 3, we see that although all protocols correlate similarly with MQMWMT, our protocols obtain the highest $\tau$ for both runs. The segment-level correlation also confirms that relying only on the error spans is not optimal and ESAspans obtains lower Kendall score.
@@
Table 4: Similarity between spans of different annotation protocols (one campaign) computed as percentage how much of $B$ does $A$ contain. For example, 93% of ESA spans were also in MQM. Any span that overlaps with another one is considered a hit, even if only with as single character.
@@ 6.4 Quality of annotations
Intra annotator agreement.
@@
Lastly, evaluating inter annotator agreement is heavily affected by the strategy of annotators, where different annotation strategy does not mean different performance of the task as Riley et al. (2024) showed. Secondly, the MQMWMT was collected with a different tooling and the documents have been presented to annotators in different order, which could also impact the inter annotator agreement.
@@
Inter annotator agreement.
@@
Agreement on the error span.
@@ 6.5 Annotation time
One of the reasons behind development of the new protocol is reducing the time requirements of the evaluation. In this section, we analyze times on our experiments only, therefore our rerun of MQM in contrast to ESA. We do not include the times for DA+SQM and MQMWMT because they were done independently outside of our study and the time data are not available. Assessing the speed of annotations is challenging as annotators took breaks during the annotation (from short breaks taking several minutes up to several hours). This makes the evaluation of time problematic and we therefore investigate the time estimate in several ways. Counting all annotators together, the median time for annotation per single paragraph for MQM is 38 seconds and for ESA is 29 seconds, a reduction of 23%. However, as median time for each annotator fluctuates, we look at the average median time across annotators. For MQM, the median is 49 seconds and for ESA it is 34 seconds, a reduction of 32%. This can be contributed mainly to the less demanding error span annotation approach used in ESA.
We note that further speedups could be made by instructing the ESA annotators to not spend extra time e.g. marking multiple error-span annotations of a single grammatical phenomenon.
@@ 7 Conclusion
A possible limitation in contrast to MQM is that now the system evaluation does not provide breakdown of error types, which could help practitioners in improving their systems. ESA does not provide this because the goal is evaluation and not diagnosis. Because of this and the costs of scaling expert labor, we are convinced that this is not a true shortcoming of ESA. Furthermore, the annotated errors can be further classified and analysed, if necessary.
@@ Appendix A User Guidelines
The following are annotation guidelines for our local ESA and MQM campaigns.
@@ A.2 MQM (Multidimensional Quality Metrics)
Figure 8: Intra annotator agreement; changes in scoring by the same annotator when evaluated again. Each point represents an annotated segment with x-axis being annotator’s score assigned in March and y-axis their score assigned in May.
@@
Appendix B Additional results
B.1 From error spans to final score
[…]
@@ B.1 From error spans to final score
B.2 Protocol evaluation withouta gold standard
Evaluating the quality of a protocol without a target to compare to is difficult. In previous sections we assumed that MQMWMT is the gold standard, which might bias the evaluation in favor of MQM. Even though the results showed higher correlations of ESA with MQMWMT. For completeness, we consider the methodology of Zouhar et al. (2024) which does not require target gold standard to compare the quality of annotation protocols.
The assumption is that annotation protocols have various levels of noise, but are unbiased in what they measure. Because MQM and DA+SQM might measure different things, we want to compare each to the perfect ranking of the particular thing they aim to measure. The linking hypothesis is that even noisy and low-quality annotation protocols would lead to the final system ranking with large enough data. Vice-versa, only robust annotation protocols would arrive at the final ranking with only a few data points per each system. This is formalized by the subset consistency accuracy. It is the system ranking accuracy on a subset of annotations with respect to the ranking induced by the full data from one annotation protocol.
We show the results in Figure 10. Out of the comparable lines, ESA (with direct scoring) achieves the highest subset consistency accuracy. In practice, this translates to needing fewer annotated examples to achieve the final system ranking. This directly corresponds to lower annotation costs.
Figure 10: Subset consistency accuracy (Zouhar et al., 2024) of annotation protocols. E.g. with just 60 annotated segments, MQMWMT achieves 95% system ranking accuracy with its final ranking based on 160 annotated segments. Values in the legend are averages, corresponding to normalized area under the curve. The only comparable lines are ESA, ESAspans, and MQMspans because they were run in the same setting with similar crowds.
Figure 11: Tutorial to ESA annotations shown at the beginning of the campaign. All tutorial segments need to be annotated correctly before continuing.